# When and How Does CLIP Enable Domain and Compositional Generalization?

### v2

@@ 5 When Does CLIP Exhibit Domain and Compositional Generalization?
Overall, these observations suggest that while domain diversity is generally beneficial for generalization, the specific relationship between the added domain(s) and the test domain plays a critical role. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 1: Domain diversity enhances domain generalization and compositional generalization. However, the relationship between domains also matters.
@@
While high diversity improves generalization, one may intuitively expect CLIP to generalize better to unseen classes within a test domain if it has seen some other classes of that domain during training. However, Figure 2 surprisingly reveals the opposite: CLIP models trained on a subset of classes from the test domain (CG low-/high-diversity) are often slightly outperformed by models that have not seen the test domain at all (Natural-only, Leave-out-domain). {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 2: Compositional generalization can be weaker than domain generalization. This finding is surprising, since the test domain is entirely unseen and domain generalization can require extrapolation. In contrast, compositional generalization is expected to perform better as it has partial exposure to that domain. However, our results suggest that this intuition may not always hold.
@@ 6.1 The Role of Visual Embeddings
As shown in Table 3, CLIP models trained on more diverse data—which generalize better—exhibit more cross-domain feature sharing in their visual embeddings. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 3: CLIP shares more features across domains in its embeddings as generalization improves.
@@ 6.2 Why Does CLIP Sometimes Fail To Generalize?
Weisfeiler-Lehman similarity ($\uparrow$) 0.218 0.268
@@
Figure 6 confirms that the Quickdraw domain noticeably differs from the other domains, both in terms of representational and mechanistic similarity, supporting our hypothesis that (compositional) generalization requires sufficient sharing of intermediate features and circuits. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 4: Sufficient sharing of intermediate features and circuits is crucial for generalization to succeed.
@@ 7 Discussion
We conducted our experiments mainly on NVIDIA RTX 2080 GPUs and estimated the total GPU hours to be approximately 25 0002500025\,00025 000. With a carbon efficiency of 0.321 kgCO2eq/kWh [^4] Estimated carbon efficiency of the German power grid in 2024 according to https://www.nowtricity.com/country/germany/, total emissions are estimated to be about 1 72517251\,7251 725 kgCO2eq, using the Machine Learning Impact calculator (Lacoste et al., 2019).
@@ A.6 CLIP Training Details
[…]
@@ C.1 Technical Details on Sparse Autoencoders
1:  $\text{hist}\leftarrow\mathbf{0}\in\mathbb{R}^{h}$ // $h$ is the number of concepts
@@
11:  $\frac{1}{|S|}\sum\limits_{s\in S}s$ // overlap averaged over domains $D_{j}$, classes $c$, and $k$
@@ C.2 Technical Details on Center Kernel Alignment and Additional Results
We computed center kernel alignment between all pairs of domains for each class. Thereby, we can assess the representational similarity for each class across the domains. For visualization, for each domain, we averaged over all classes to obtain a representational similarity estimate. In the main text, we further averaged together the non-Quickdraw domains (all domains are shown in Figures 16 and 17).
@@
Figures 16 and 17 show results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$. Interestingly, we find that the representational gap widens for the unseen classes $C_{2}$.
@@
Figure 16: Linear center kernel alignment (CKA) similarity. Quickdraw has the lowest representational similarity across domains. This particularly emphasized for the unseen classes $C_{2}$ (16(c), 16(f)).
@@
Figure 17: Non-linear center kernel alignment (CKA) similarity. The results follow the same pattern as the linear CKA (Figures 6(a) and 16): Quickdraw is the most representational dissimilar domain.
@@ C.3 Technical Details on the Circuit Similarity Analysis and Additional Results
We compared the circuit similarity of pairs of graphs resorting to classical graph similarity measures.
@@
$\frac{|N_{C_{\text{domain}~{}D_{1}\text{, class c}}}\cap N_{C_{\text{domain}~{}D_{2}\text{, class c}}}|}{|N_{C_{\text{domain}~{}D_{1}\text{, class c}}}\cup N_{C_{\text{domain}~{}D_{2}\text{, class c}}}|}\quad.$ (10)
@@
    Counting occurences: Count the occurrences of each compressed label to obtain the feature vectors $\phi_{h}(G_{1}),\;\phi_{h}(G_{2})$.
@@
For our analysis, we used $h=3$ iterations. Our implementation is based on the publicly available code from https://github.com/emanuele/jstsp2015, License MIT.
@@
Figures 18 and 13 shows results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$. Interestingly, we find that, similar as for the representational similarities in Figures 16 and 17, nodes are less shared and circuits are slightly less similar for the unseen classes $C_{2}$.
(a) All classes.
(b) Seen classes only.
(c) Unseen classes only.
(d) All classes.
(e) Seen classes only.
(f) Unseen classes only.
Figure 18: Amount of shared neurons. The Quickdraw domains shares the least neurons. This is especially apparent for the unseen classes $C_{2}$ (18(c), 18(f)).
Table 13: Weisfeiler-Lehman similarities. Higher similarities mean higher circuit (graph) similarity. The Quickdraw domain exhibits the least degree of similarity, supporting our hypothesis that sharing of the circuits is critical for generalization.
Clipart Infograph Painting Quickdraw Real Sketch
All classes 0.281 0.246 0.266 0.218 0.273 0.273
Seen classes only 0.281 0.246 0.265 0.218 0.273 0.273
Unseen classes only 0.278 0.258 0.274 0.211 0.272 0.275

### v3

@@ 5 When Does CLIP Exhibit Domain and Compositional Generalization?
Overall, these observations suggest that while domain diversity is generally beneficial for generalization, the specific relationship between the added domain(s) and the test domain plays a critical role. Finding 1: Domain diversity enhances domain generalization and compositional generalization. However, the relationship between domains also matters.
@@
While high diversity improves generalization, one may intuitively expect CLIP to generalize better to unseen classes within a test domain if it has seen some other classes of that domain during training. However, Figure 2 surprisingly reveals the opposite: CLIP models trained on a subset of classes from the test domain (CG low-/high-diversity) are often slightly outperformed by models that have not seen the test domain at all (Natural-only, Leave-out-domain). Finding 2: Compositional generalization can be weaker than domain generalization. This finding is surprising, since the test domain is entirely unseen and domain generalization can require extrapolation. In contrast, compositional generalization is expected to perform better as it has partial exposure to that domain. However, our results suggest that this intuition may not always hold.
@@ 6.1 The Role of Visual Embeddings
As shown in Table 3, CLIP models trained on more diverse data—which generalize better—exhibit more cross-domain feature sharing in their visual embeddings. Finding 3: CLIP shares more features across domains in its embeddings as generalization improves.
@@ 6.2 Why Does CLIP Sometimes Fail To Generalize?
Weisfeiler-Lehman similarity ($\uparrow$) 0.218 0.28
@@
Figure 6 confirms that the Quickdraw domain noticeably differs from the other domains, both in terms of representational and mechanistic similarity, supporting our hypothesis that (compositional) generalization requires sufficient sharing of intermediate features and circuits. Finding 4: Sufficient sharing of intermediate features and circuits is crucial for generalization to succeed.
@@ 7 Discussion
We conducted our experiments mainly on NVIDIA RTX 2080 GPUs and estimated the total GPU hours to be approximately 25 00025\,000. With a carbon efficiency of 0.321 kgCO2eq/kWh [^4] Estimated carbon efficiency of the German power grid in 2024 according to https://www.nowtricity.com/country/germany/., total emissions are estimated to be about 1 7251\,725 kgCO2eq, using the Machine Learning Impact calculator (Lacoste et al., 2019).
@@ A.6 CLIP Training Details
Following Fang et al. (2022), we trained CLIP models with an embedding size of 1024 with ResNet-50 (He et al., 2016) and a transformer text encoder (Vaswani et al., 2017) (12 layers with a width of 512, 8 attention heads, and context length of 77). We trained the models for 32 epochs with a batch size of 1024 with AdamW (learning rate of 0.001, $\beta_{1}=\text{0.9}$, $\beta_{2}=\text{0.999}$, $\epsilon=\text{1e-8}$, weight decay of 0.2) and cosine annealing learning rate scheduling with 500 warmup steps. We used the code from OpenCLIP (Cherti et al., 2023) (https://github.com/mlfoundations/open_clip, License: custom) for our training implementation, including OpenCLIP’s default augmentations.
@@ C.1 Technical Details on Sparse Autoencoders
[…]
@@
[…]
@@ C.2 Technical Details on Center Kernel Alignment and Additional Results
We computed center kernel alignment between all pairs of domains for each class. Thereby, we can assess the representational similarity for each class across the domains. For visualization, for each domain, we averaged over all classes to obtain a representational similarity estimate. In the main text, we further averaged together the estimates for the non-Quickdraw domains (‘Others’). The estimates for each individual domain is shown in Figures 16 and 17.
@@
Figure 16: Linear center kernel alignment (CKA) similarity. Quickdraw has the lowest representational similarity across domains. This particularly emphasized for the unseen classes $C_{2}$ (16(c), 16(f)). We used the CLIP model trained on only domain-invariant captions from Figures 5(b) and 4 (second row). Figure 18 provides results for other CLIP models.
@@
Figure 17: Non-linear center kernel alignment (CKA) similarity. The results follow the same pattern as the linear CKA (Figures 6(a) and 16): Quickdraw is the most representational dissimilar domain. We used the CLIP model trained on only domain-invariant captions from Figures 5(b) and 4 (second row).
(a) Clipart.
(b) Infograph.
(c) Painting.
(d) Quickdraw.
(e) Sketch.
Figure 18: Linear center kernel alignment (CKA) similarity for different CLIP models. We used the CLIP models from the CG high-diversity setting in Figure 2 and report representational similarities on the unseen classes $C_{2}$. The patterns for all classes $C$ and for the seen classes $C_{1}$ are qualitatively similar. Among the domains, Clipart, Painting, and Sketch exhibit comparable levels of representational similarity, whereas Infograph and particularly Quickdraw show lower similarities.
Figures 16 and 17 show results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$ for the CLIP model solely trained on domain-invariant captions from Figures 5(b) and 4 (second row). Interestingly, we find that the representational gap widens for the unseen classes $C_{2}$. Figure 18 shows the representational similarity for different CLIP models from Figure 2 for the unseen classes $C_{2}$. Similar to the above, the CLIP model evaluated on the Quickdraw domain exhibits low representational similarity to models from other domains. In contrast, the CLIP models with the test domains Clipart, Infograph, Painting, and Sketch show higher or more comparable representational similarities.
@@ C.3 Technical Details on the Circuit Similarity Analysis and Additional Results
(a) All classes.
(b) Seen classes only.
(c) Unseen classes only.
(d) All classes.
(e) Seen classes only.
(f) Unseen classes only.
Figure 19: Amount of shared neurons. The Quickdraw domains shares the least neurons. This is especially apparent for the unseen classes $C_{2}$ (19(c), 19(f)). Results are for the CLIP model trained on only domain-invariant captions from Figures 5(b) and 4 (second row). Figure 20 provides results for other CLIP models.
Table 13: Weisfeiler-Lehman similarities. Higher similarities mean higher circuit (graph) similarity. We used the CLIP model trained on only domain-invariant captions from Figures 5(b) and 4 (second row); results for other CLIP models are provided in Table 14. The Quickdraw domain exhibits the least degree of similarity, supporting our hypothesis that sharing of circuits is critical for generalization.
Clipart Infograph Painting Quickdraw Real Sketch
All classes 0.281 0.246 0.266 0.218 0.273 0.273
Seen classes only 0.281 0.246 0.265 0.218 0.273 0.273
Unseen classes only 0.278 0.258 0.274 0.211 0.272 0.275
(a) Clipart.
(b) Infograph.
(c) Painting.
(d) Quickdraw.
(e) Sketch.
Figure 20: Amount of shared neurons for different CLIP models. We used the CLIP models from the CG high-diversity setting in Figure 2 and report results on the unseen classes $C_{2}$. While Clipart, Infograph, Painting, and Sketch exhibit similar levels of neuron sharing, Quickdraw shows significantly less sharing.
Table 14: Weisfeiler-Lehman similarities for different CLIP models. Higher similarities mean higher circuit (graph) similarity. We used the CLIP models in the CG high-diversity setting from Figure 2. Results are for the unseen classes $C_{2}$. The CLIP models with the test domains Clipart, Painting, and Sketch have a higher degree of similarity than Infograph and Quickdraw.
Test domain Test domain vs. Others Others vs. Others
Clipart 0.274 0.243
Infograph 0.246 0.250
Painting 0.263 0.248
Quickdraw 0.176 0.272
Sketch 0.259 0.242
@@
We compared the circuit similarity of pairs of circuits (graphs) resorting to classical graph similarity measures. Similar to the representational analysis, we computed the graph similarity measures between all pairs of domains for each class and averaged over all classes to obtain the final circuit similarity estimates. In the main text, we further averaged together the estimates of the non-Quickdraw domains (‘Others’). The estimates of each individual domain are shown in Figures 19(d), 19(e), 19(f) and 13.
@@
[…]
@@
    Counting occurrences: Count the occurrences of each compressed label to obtain the feature vectors $\phi_{h}(G_{1}),\;\phi_{h}(G_{2})$.
@@
For our analysis, we used $h=3$ iterations. Our implementation is based on the publicly available code from https://github.com/emanuele/jstsp2015, License: MIT.
@@
Figures 19 and 13 show results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$ for the CLIP model only trained on domain-invariant captions from Figures 5(b) and 4 (second row). Neurons (nodes) are less shared and the circuits are slightly less similar for the unseen classes $C_{2}$. Figure 20 and Table 14 extend this analysis to the different CLIP models from Figure 2 (results are on the unseen classes $C_{2}$). Across these models, we consistently find that domains like Clipart, Painting, and Sketch maintain similar levels of neuron and circuit sharing, whereas Quickdraw exhibits less sharing. Figure 21 shows that our observations are robust across various neuron thresholds (we chose to retain the 10% most important neurons per layer in Figures 6(b), 6(c), 19, 20, 13 and 14).
(a) Clipart.
(b) Infograph.
(c) Painting.
(d) Quickdraw.
(e) Sketch.
Figure 21: Amount of shared neurons for different CLIP models over various neuron thresholds. We used the CLIP models from the CG high-diversity setting in Figure 2 and report results on the unseen classes $C_{2}$, averaged across all layers and evaluated over multiple thresholds. The patterns observed in Figure 20 remain consistent across these thresholds.